im trying to figure out how to get k from this equation: T= 2pi times the square root of m/k. I want to know how to rearrange this in order to get k.
Answer by stanbon(48564)

figure out how to get k from this equation: T= 2pi times the square root of m/k.
--------------
T = 2(pi)sqrt(m/k)
sqrt(m/k) = t/(2pi)
Square both sides to get:
m/k = (t/2pi)^2
Invert both sides to get:
k/m = (2pi/t)^2
Multiply both sides by m to get:
k = m(2pi/t)^2
===============
Cheers,
Stan H.